OpenStudy (anonymous):

What would be the total cost of 6 ink cartridges at $20 each and 2 reams of paper at $7.00 each? Taxes are 7 cents on the dollar (7%).

OpenStudy (texaschic101):

6 ink cartridges at $ 20 each = (6)(20) = $120 2 reams of paper at $7 each = (2)(7) = $14 120 + 14 = 134 (without taxes) The sales tax is 7%, turned to a decimal = .07 Multiply that by the cost without taxes and that is your sales tax, then add the sales tax to the cost before taxes and you will have your total cost, tax and all. 134(.07) + 134 = cost with taxes 9.38 + 134 = $143.38

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No. Thank you very much! :)

OpenStudy (texaschic101):

no problem :)